# Extremum Seeking with High-Order Lie Bracket Approximations: Achieving Exponential Decay Rate

### v2

@@ I Introduction
$\dot{x}=f_{0}(x)+\sum\limits_{i=1}^{m}\frac{1}{\varepsilon^{p_{i}}}f_{i}(x)u_{i}\left(k_{i}\frac{t}{\varepsilon}\right),\$ (1)
where $x\in\mathbb{R}^{n}$ is the state space vector, $p_{i}\in(0,1)$, $0<\varepsilon\ll 1$, $f_{0}$ is the drift (uncontrolled) vector field of the system, $f_{i}$ are the control vector fields, $u_{i}$ are the control inputs, $m\in Z^{+}$ is the number of control inputs, and $k_{i}\in Q_{>0}$ is a positive rational number. Control-affine systems of the form (1) characterize many real-world systems and applications, such as but not limited to robotic, multi-agent and flight dynamic systems (e.g., [1, 2, 3, 4, 5]). The geometric nature of systems of the form (1) allows the application of geometric control methods and techniques [6, 7] that are based on Lie-bracket-based approximations and analysis [8, 9, 10, 11, 12, 13, 14, 15]. A Lie bracket variation between the vector fields $f_{i}\text{ and }f_{j}$ is defined as:
$[f_{i},f_{j}]:=\frac{\partial f_{j}}{\partial x}f_{i}-\frac{\partial f_{i}}{\partial x}f_{j}.$ (2)
Lie-bracket-based approximations and analysis have been used to study motion planning of underactuated systems, including with nonholonomic constraints (e.g., [6, 10, 9, 16, 17, 18, 19, 20]) and model-free optimization and control via extremum seeking (ES) approaches [11, 12, 13, 14, 15]. In this paper, we focus on Lie bracket approximations relevant to ES systems.
Extremum seeking (ES) systems [21] are model-free and real-time dynamic optimization and control techniques that aim at stabilizing a dynamical system about the extremum point of an objective function, for which we have access to its measurements but not its expression; see [22] for a comprehensive review. ES methods referred to as classic-based (e.g., [23, 24, 25]) rely on classical averaging methods [26, 27] for analysis and design. On the other hand, control-affine ES methods [11, 12, 13, 14, 15] (the focus of this paper) rely on Lie-bracket approximations for analysis and design.
Now we briefly summarize the development of Lie bracket approximations of control-affine ES systems. For a particular case of (1) when $p_{i}=0.5,k_{i}=1\text{ for all }i$, a first-order Lie bracket system (LBS) approximation was given in [11], which also admits drift vector field and is defined not only for underactuated systems. In [28], the authors utilized the first-order LBS in [11] and provided a universal semi-global stabilizing control law based on Lyapunov function candidates; they also [29] extended the first-order LBS in [11] to a non-smooth setting. Different control law from [11] with bounded update rate was provided in [12] using first-order LBS approximation. In [13], a generalized formulation and first-order LBS approach was provided for a class of control-affine ES systems that: (i) unifies previous works such as [11, 12], and (ii) provides a new control law with enhanced stability guarantees. It is important to emphasize that ES methods based on first-order LBS approximation (e.g., [11, 12, 13]) are gradient-based ES approaches; that is, the ES system is designed such that its LBS/average behaves as a gradient-flow. In [14], the authors generalized further the conditions of $p_{i}$ and $k_{i}$ compared to [11], and introduced a second-order LBS, which approximates (1). This enabled the introduction of a Newton-based ES approach because the inclusion of second-order Lie brackets provided access to second-order derivative information (i.e., Hessian), hence enabling a Newton-based flow. Recently, the authors in [15] provided a generalized approach for generating LBSs with higher-orders to approximate (1). They also showed that LBS approximations are averaging terms themselves, which guarantees the closeness of trajectories between LBSs and the original system (1) given small enough $\varepsilon$. Moreover, it was observed that ES designs based on third-order LBS possess a faster convergence rate when compared with even Newton-based ES; the authors in [15] argued (without proof) that the observed faster convergence is due to the inclusion of third-order Lie brackets, which provide access to higher-order derivatives beyond Hessian (i.e., third-order derivative information).
@@
Motivation & contributions. Inspired by the concept of higher-order Lie bracket averaging, this paper further explores the application of these techniques to extremum seeking problems, with a particular focus on achieving faster convergence rates. In particular, many extremum seeking algorithms based on first-order Lie bracket approximations exhibit exponential convergence when the cost function $J(x)$ behaves locally like a quadratic function near the extremum point $x^{*}$ (i.e., $J(x)\sim\|x-x^{*}\|^{2}$). However, if the cost function behaves like a higher-degree polynomial near the extremum, i.e., $J(x)\sim\|x-x^{*}\|^{m}$ with $m>2$, such algorithms exhibit only a polynomial decay rate. Even though it was observed in [15] that using higher-order Lie brackets to design ES may lead to faster convergence rate when $J(x)\sim\|x-x^{*}\|^{4}$, no conclusion or a proof was provided regarding the nature of the faster convergence resulting from higher-order Lie brackets.
@@
In this paper, we investigate a class of extremum seeking problems involving the unconstrained minimization of a cost function $J(x)$. The function $J$ is unknown in terms of an explicit analytic expression, but it can be evaluated (measured) at any point. We focus on designing an ES control system of the form $\dot{x}=u(t,J(x)),$ such that the system’s trajectories tend asymptotically to an extremum point of the function $J$. To develop our approach, in this paper we assume that the cost function behaves locally like an $m$-th order polynomial near the minimizer $x^{*}\in\mathbb{R}^{n}$, i.e., $J(x)\sim\|x-x^{*}\|^{m}$, with some $m\geq 2$.
@@
The main contribution of this paper is a novel extremum seeking design framework that leverages the excitation of higher-order Lie brackets to steer the system along directions corresponding to higher-order derivatives of the cost function. This approach helps to increase the convergence rate, in particular, ensuring the exponential convergence even for cost functions that are not quadratic in nature. Furthermore, we generalize the result of [13], which described a family of vector fields whose first-order Lie bracket equals the gradient of the cost function. In this paper, we extend this idea by deriving a formula that generates vector fields such that the corresponding $\ell$-th order Lie bracket equals the $\ell$-th derivative of the cost function. In addition, we discuss several strategies for exciting Lie brackets using different numbers of dither signals, which gives higher flexibility in the design of control vector fields.
The remainder of the paper is organized as follows. Section II introduces main notations and definitions and preliminary results related to Lie bracket-based extremum seeking control, and describes the main idea behind our approach. Section III presents the main contributions of this paper, which are illustrated through numerical simulations in Section IV. Finally, Section V provides concluding remarks and outlines potential extensions and applications of the proposed framework.
@@ II-A Notations
Throughout the text, $\mathbb{R}^{+}=[0,\infty)$ denotes the set of all non-negative real numbers;
$B_{\delta}(x^{*})$ is the $\delta$-neighborhood of $x^{*}{\in}\mathbb{R}^{n}$, $\overline{B_{\delta}(x^{*})}$ is its closure;
for $h{\in}C^{1}(\mathbb{R}^{n};\mathbb{R})$, $\xi{\in}\mathbb{R}^{n}$, we denote $\nabla h(\xi):=\frac{\partial h(x)}{\partial x}^{T}\Big{|}_{x=\xi}$ to be a column vector;
for $h\in C^{N}(\mathbb{R};\mathbb{R})$, we define its $\ell$-th order derivative as $h^{(\ell)}(x):=\frac{d^{\ell}h(x)}{dx^{\ell}}$, for any $\ell\in\{0,\dots,N\}$, with $h^{(0)}(x):=h(x)$;
for a function $f:\mathbb{R}\to\mathbb{R}$, $f(z)=O(z)$ as $z\to 0$ means that there is a $c>0$ such that $|f(z)|\leq c|z|$ in some neighborhood of 00;
for $f,g:\mathbb{R}^{n}\to\mathbb{R}^{n}$, $x\in\mathbb{R}^{n}$, the Lie derivative is defined as $L_{g}f(x)=\lim\limits_{s\to 0}\dfrac{f(x+sg(x))-f(x)}{s}$,
and $[f,g](x)=L_{f}g(x)-L_{g}f(x)$ is the first order Lie bracket.
to define higher-order Lie brackets, we introduce $\ell$-dimensional multi-index $I_{\ell}=(i_{1},\dots,i_{\ell})$; then
$f_{I_{\ell}}(x)=\left[\big{[}[f_{i_{1}},f_{i_{2}}],f_{i_{3}}\big{]},\dots,f_{i_{\ell}}\right](x)$ – the right-iterated Lie bracket of length $\ell$, or $(\ell-1)$-order Lie bracket; for $I_{1}{\in}\{1,...,m\}$, $f_{I_{1}}$ simply means a corresponding vector field;
$L^{\infty}_{[0,\varepsilon]}$ – the class of essentially bounded measurable functions on $[0,\varepsilon]$.
II-B Lie brackets approximations
[…]
@@ II-B Lie brackets approximations
with
$L_{1}$ $=0,$
$L_{2}$ $=\sum_{j_{1}=1}^{m}\sum_{j_{2}=j_{1}+1}^{m}\nu_{j_{1}j_{2}}[f_{j_{1}},f_{j_{2}}],$
$L_{3}$ $=\sum_{j_{1}=1}^{m}\sum_{j_{2}=j_{1}+1}^{m}\sum_{j_{3}=1}^{m}\nu_{j_{1}j_{2}j_{3}}[f_{j_{3}},[f_{j_{1}},f_{j_{2}}]],$
$\begin{split}L_{4}&=\sum_{j_{1}=1}^{m}\sum_{j_{2}=j_{1}+1}^{m}\sum_{j_{3}=1}^{m}\sum_{j_{4}=j_{3}+1}^{m}\beta_{1_{j_{1}j_{2}j_{3}j_{4}}}\Big{[}[f_{j_{1}},f_{j_{2}}],[f_{j_{3}},f_{j_{4}}]\Big{]}\\ &+\sum_{j_{1}=1}^{m}\sum_{j_{2}=j_{1}+1}^{m}\sum_{j_{3}=1}^{m}\sum_{j_{4}=1}^{m}\beta_{2_{j_{1}j_{2}j_{3}j_{4}}}[[[f_{j_{1}},f_{j_{2}}],f_{j_{3}}],f_{j_{4}}],\end{split}$
where $\nu_{j_{1}j_{2}}$, $\nu_{j_{1}j_{2}j_{3}}$, $\beta_{1_{j_{1}j_{2}j_{3}j_{4}}}$ and $\beta_{2_{j_{1}j_{2}j_{3}j_{4}}}$ are coefficients resulting from the iterated integrals of the dither input signals (formulas are provided in [15, Section 4]). Truncating (3) at $r=2$ provides first-order LBS (e.g., [11, 13]). Similarly, truncating (3) at $r=3$ and $r=4$ provides second- and third-order LBS, respectively. The stability properties of systems (1) and (3) are related as follows.
@@
Below we recall the notion of practical stability.
Definition 1
A compact set $S\subset\mathbb{R}^{n}$ is said to be locally practically uniformly asymptotically stable for (1) if:
– it is practically uniformly stable, i.e. for every $\rho>0$ there exist $\delta>0$ and $\bar{\varepsilon}>0$ such that, for all $t_{0}\geq 0$ and $\varepsilon\in(0,\bar{\varepsilon})$, any solution of (1) with $x^{0}\in B_{\delta}(S)$ satisfies $x(t)\in B_{\varepsilon}(S)$ for all $t\geq t_{0}$;
$\hat{\delta}$-practically uniformly attractive with some $\hat{\delta}>0$, i.e. for every $\rho>0$ there exist $t_{\rho}\geq 0$ and $\bar{\varepsilon}>0$ such that, for all $t_{0}\geq 0$ and $\varepsilon\in(0,\bar{\varepsilon})$, any solution of (1) with $x^{0}\in B_{\delta}(S)$ satisfies $x(t)\in B_{\varepsilon}(S)$ for all $t\geq t_{0}+t_{1}$.
If the attractivity property holds for every $\hat{\delta}{>}0$, then the set $S$ is called semi-globally practically uniformly asymptotically stable for (1).
@@
$\dot{x}=\frac{1}{\sqrt{\varepsilon}}\Big{(}F_{1}(J(x))u_{1}^{\varepsilon}(t)+F_{2}(J(x))u_{2}^{\varepsilon}(t)\Big{)},$ (4)
where $F_{1}\circ J,F_{2}\circ J\in C^{2}(D)$ satisfy the relation $[F_{1},F_{2}](Z)=1$ for all $z\in\mathbb{R}$ (which implies $[F_{1}\circ J,F_{2}\circ J](x)=\nabla J(x)$), $u_{1}^{\varepsilon}(t)=2\sqrt{{\pi}{\varepsilon^{-1}}}\cos\big{(}{2\pi t}{\varepsilon^{-1}}\big{)}$, $u_{2}^{\varepsilon}(t)=2\sqrt{{\pi}{\varepsilon^{-1}}}\sin\big{(}{2\pi t}{\varepsilon^{-1}}\big{)}$, and assume that the cost function $J$ satisfies the following properties in some domain $D\subseteq\mathbb{R}^{n}$:
@@
    there exists an $x^{*}{\in}D$ such that
    $\nabla J(x)=0\text{ if and only if }x=x^{*},$
    and $J(x^{*})=J^{*}{\in}\mathbb{R}$, $J(x)>J(x^{*})$ for all $x\in D{\setminus}\{x^{*}\}$.
@@
    There exist constants $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2},\mu$, and $m\geq 1$, such that, for all $x\in D$,
@@
    $\left\|\frac{\partial^{2}J(x)}{\partial x^{2}}\right\|{\leq}$ $\mu(J(x)-J^{*})^{1{-}\frac{2}{m}}.$
@@
It can be shown that the point $x=x^{*}$ is practically asymptotically stable for (4), with the convergence rate dependent on the parameter $m$ in A1.2). Namely, the following result follows from [13]:
@@
    $\|x(t)-x^{*}\|\leq\Big{(}c_{m}\|x^{0}-x^{*}\|^{2-m}+\tilde{c}_{m}t\Big{)}^{-1/(m-2)}+\rho.$
Here $c_{m},\tilde{c}_{m}$ are positive constants; $c_{m}$ can be made arbitrarily close to 1 by choosing a sufficiently small $\varepsilon$.
More technical details on the above decay rate estimates can be found in [13].
For the sake of clarity, we have assumed here $x\in\mathbb{R}$, however, the above result can be easily extended to the case $x\in\mathbb{R}^{n}$ (see, e.g., [11, 13]).

### v3

@@ I Introduction
$\dot{x}=f_{0}(x)+\sum_{i=1}^{m}{\varepsilon^{-p_{i}}}f_{i}(x)u_{i}\left(k_{i}{t}/{\varepsilon}\right),\\$ (1)
where $x\in\mathbb{R}^{n}$ is the state space vector, $p_{i}\in(0,1)$, $0<\varepsilon\ll 1$, $f_{0}$ is the drift (uncontrolled) vector field of the system, $f_{i}$ are the control vector fields, $u_{i}$ are the control inputs, $m\in\mathbb{N}$ is the number of control inputs, and $k_{i}\in\mathbb{Q}$. Control-affine systems of the form (1) arise in many real-world applications, including (but not limited to) robotics, multi-agent systems, and flight dynamics (e.g., [1, 2, 3, 4]). The nature of system (1) allows applications of geometric control methods and techniques relying on Lie bracket approximation ideas[5, 6, 7, 8, 9, 10, 11, 12, 13]. Such approaches have been widely used in the study of motion planning for underactuated systems, including those with nonholonomic constraints (e.g., [7, 8, 6, 14, 15, 16, 17, 18]), as well as in model-free optimization and control via extremum seeking (ES) approaches [9, 10, 11, 12, 13], which is the main focus of this paper. Extremum seeking is model-free, real-time dynamic optimization and control framework aimed at stabilizing a dynamical system about the extremum point of an objective function that is accessible through measurements but whose analytical form is unknown [19, 20]. ES methods referred to as classic-based (e.g., [21, 22, 23]) rely on classical averaging methods [24, 25] for analysis and design. Another family of approaches exploit control-affine ES systems and rely on Lie bracket approximation for analysis and design, such as, e.g., [9, 10, 11, 12, 13]. In particular, for the case $p_{i}=0.5,k_{i}=1\text{ for all }i$, a first-order Lie bracket system (LBS) approximation of (1) and ES control laws were given, e.g., in [9, 10, 11]. Note that such methods define an ES system whose first-order LBS behaves like a gradient flow of the objective function. In [12], the conditions on $p_{i}$ and $k_{i}$ were further generalized compared to [9], and a second-order LBS was introduced to approximate (1). This extension enabled the introduction of a Newton-based ES approach, since the inclusion of second-order Lie brackets provided access to second-order derivative information (i.e., to the Hessian), and thus allowed to have a behavior similar to a Newton-type flow. In the recent paper [13], the authors proposed a generalized framework for constructing higher-order LBSs to approximate (1). They also showed that LBS approximations themselves are averaging terms, which guarantees the closeness of trajectories between LBSs and the original system (1), provided that $\varepsilon$ is small enough. Moreover, it was observed that ES designs based on third-order LBS possess a faster convergence rate when compared with even Newton-based ES; the authors in [13] argued (without proof) that the observed faster convergence is due to the inclusion of third-order Lie brackets, which provide access to higher-order derivatives beyond Hessian (i.e., third-order derivative information).
@@
Motivation & contributions. Inspired by the concept of higher-order Lie bracket averaging, this paper further explores the application of these techniques to ES problems, with a particular focus on achieving faster convergence rates. In particular, many ES algorithms based on first-order Lie bracket approximation exhibit exponential convergence when the cost function $J(x)$ behaves locally like a quadratic function near the extremum point $x^{*}$ (i.e., $J(x)\sim\|x-x^{*}\|^{2}$). However, if the cost function behaves like a higher-degree polynomial near the extremum, i.e., $J(x)\sim\|x-x^{*}\|^{m}$ with $m>2$, such algorithms exhibit only a polynomial decay rate [11]. Even though it was observed in [13] that using higher-order Lie brackets to design ES may lead to faster convergence rate when $J(x)\sim\|x-x^{*}\|^{4}$, no conclusion or a proof was provided regarding the nature of the faster convergence resulting from higher-order Lie brackets.
@@
In this paper, we provide a preliminary study of a class of ES problems related to the unconstrained minimization of a cost function $J(x)$. The function $J$ is unknown in terms of an explicit analytic expression, but it can be evaluated (measured) at any point. We focus on designing an ES control system of the form $\dot{x}=u(t,J(x)),$ such that the system’s trajectories tend to an extremum point of $J$. To develop our approach, we assume that the cost function behaves locally like an $m$-th order polynomial near the minimizer $x^{*}\in\mathbb{R}^{n}$, i.e., $J(x)\sim\|x-x^{*}\|^{m}$, with some $m\geq 2$.
@@
The main contribution of this paper is a novel ES design framework that leverages the excitation of higher-order Lie brackets to steer the system along directions corresponding to higher-order derivatives of the cost function. This approach helps to increase the convergence rate, in particular, ensuring the exponential convergence even for cost functions that are not quadratic in nature. Furthermore, we generalize the result of [11], which described a family of vector fields whose first-order Lie bracket equals the gradient of the cost function. In this paper, we extend this idea by deriving a formula that generates vector fields such that the corresponding $\ell$-th order Lie bracket equals the $\ell$-th derivative of the cost function. In addition, we discuss several strategies for exciting Lie brackets using different numbers of dither signals, which gives higher flexibility in the design of control vector fields.
@@ II-A Notations
Throughout the text, $\mathbb{R}^{+}=[0,\infty)$ denotes the set of all non-negative real numbers; $B_{\delta}(x^{*})$ is the $\delta$-neighborhood of $x^{*}{\in}\mathbb{R}^{n}$, $\overline{B_{\delta}(x^{*})}$ is its closure; for $h{\in}C^{1}(\mathbb{R}^{n};\mathbb{R})$, $\xi{\in}\mathbb{R}^{n}$, we denote $\nabla h(\xi):=\frac{\partial h(x)}{\partial x}^{T}\Big|_{x=\xi}$ to be a column vector; for $h\in C^{N}(\mathbb{R};\mathbb{R})$, we define its $\ell$-th order derivative as $h^{(\ell)}(x):=\frac{d^{\ell}h(x)}{dx^{\ell}}$, for any $\ell\in\{0,\dots,N\}$, with $h^{(0)}(x):=h(x)$; for an $f:\mathbb{R}\to\mathbb{R}$, $f(z)=O(z)$ as $z\to 0$ means that there is a $c>0$ such that $|f(z)|\leq c|z|$ in a neighborhood of 0; for $f,g:\mathbb{R}^{n}\to\mathbb{R}^{n}$, $x\in\mathbb{R}^{n}$, the Lie derivative is defined as $L_{g}f(x)=\lim_{s\to 0}\frac{f(x+sg(x))-f(x)}{s}$, and $f,g](x)=L_{f}g(x)-L_{g}f(x)$ is the first order Lie bracket. To define higher-order Lie brackets, we introduce $\ell$-dimensional multi-index $I_{\ell}=(i_{1},\dots,i_{\ell})$; then $f_{I_{\ell}}(x)=\left[\big[[f_{i_{1}},f_{i_{2}}],f_{i_{3}}\big],\dots,f_{i_{\ell}}\right](x)$ – the right-iterated Lie bracket of length $\ell$, or $(\ell-1)$-order Lie bracket; for $I_{1}{\in}\{1,...,m\}$, $f_{I_{1}}$ denotes a corresponding vector field; $L^{\infty}_{[0,\varepsilon]}$ – the class of essentially bounded measurable functions on $0,\varepsilon]$.
II-B Lie bracket approximations
Consider the control-affine system (1), where $x=(x_{1},\dots,x_{n})^{T}{\in}\mathbb{R}^{n}$ is the state vector, $x(t_{0})=x^{0}{\in}\mathbb{R}^{n}$ (without loss of generality, we assume $t_{0}=0$), $\varepsilon>0$ is a small parameter, $f_{i}{:\mathbb{R}^{n}\to\mathbb{R}^{n}}$ are continuously differentiable (up to any order) vector fields , and $u_{i}$ are continuous in $t$ and $T$-periodic functions with zero average, i.e., $\int_{0}^{T}u_{i}(\tau)d\tau=0$. LBS approximations of (1) up to a third-order are [13]:
@@ II-B Lie bracket approximations
with $L_{1}=0$, $L_{2}=\sum_{\underset{j_{2}=j_{1}+1}{j_{1}=1}}^{m}[f_{j_{1}},f_{j_{2}}]$,
$L_{3}=\sum_{\underset{j_{2}=j_{1}+1}{j_{1},j_{3}=1}}^{m}\nu_{j_{1}j_{2}j_{3}}[f_{j_{3}},[f_{j_{1}},f_{j_{2}}]]$,
$L_{4}=\sum_{{\underset{j_{2}=j_{1}+1}{j_{1},j_{3}=1}}}^{m}\big(\sum_{j_{4}=1}^{m}\beta_{1_{j_{1}j_{2}j_{3}j_{4}}}\Big[[f_{j_{1}},f_{j_{2}}],[f_{j_{3}},f_{j_{4}}]\Big]+\sum_{j_{4}=j_{3}+1}^{m}\beta_{2_{j_{1}j_{2}j_{3}j_{4}}}[[[f_{j_{1}},f_{j_{2}}],f_{j_{3}}],f_{j_{4}}]\big)$, where $\nu_{j_{1}j_{2}}$, $\nu_{j_{1}j_{2}j_{3}}$, $\beta_{1_{j_{1}j_{2}j_{3}j_{4}}}$ and $\beta_{2_{j_{1}j_{2}j_{3}j_{4}}}$ are coefficients resulting from the iterated integrals of the dither input signals (formulas are provided in [13, Section 4]). Truncating (2) at $r=2$ provides first-order LBS (e.g., [9, 11]). Similarly, truncating (2) at $r=3$ and $r=4$ provides second- and third-order LBS, respectively. The stability properties of systems (1) and (2) are related as follows.
@@
We omit the notions of practical asymptotic stability and practical exponential stability because of the space limits and refer the readers to, e.g., [9, 11].
@@
$\dot{x}=\varepsilon^{-1/2}\Big(F_{1}(J(x))u_{1}^{\varepsilon}(t)+F_{2}(J(x))u_{2}^{\varepsilon}(t)\Big),$ (3)
where $F_{1}\circ J,F_{2}\circ J\in C^{2}(D)$ satisfy $F_{1},F_{2}](Z)=1$ for all $z\in\mathbb{R}$ (which implies $F_{1}\circ J,F_{2}\circ J](x)=\nabla J(x)$), $u_{1}^{\varepsilon}(t)=2\sqrt{{\pi}{\varepsilon^{-1}}}\cos\big({2\pi t}{\varepsilon^{-1}}\big)$, $u_{2}^{\varepsilon}(t)=2\sqrt{{\pi}{\varepsilon^{-1}}}\sin\big({2\pi t}{\varepsilon^{-1}}\big)$, and assume that the cost function $J$ satisfies the following properties in some domain $D\subseteq\mathbb{R}^{n}$:
@@
    there is an $x^{*}{\in}D$ s. t. $\nabla J(x){=}0\text{ if and only if }x=x^{*},$ and $J(x^{*})=J^{*}{\in}\mathbb{R}$, $J(x)>J(x^{*})$ for all $x\in D{\setminus}\{x^{*}\}$;
@@
    there are $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2},\mu$, and $m\geq 1$, such that
@@
    $\left\|\frac{\partial^{2}J(x)}{\partial x^{2}}\right\|{\leq}$ $\mu(J(x)-J^{*})^{1{-}\frac{2}{m}}\text{ for all }x\in D.$
@@
As follows from [11], under the above conditions the point $x=x^{*}$ is practically asymptotically stable for (3), with the convergence rate dependent on the parameter $m$ in A1.2):
@@
    $\|x(t)-x^{*}\|\leq\Big(c_{m1}\|x^{0}-x^{*}\|^{2-m}+c_{m2}t\Big)^{-1/(m-2)}+\rho,$ with some $c_{m},c_{m1},c_{m2}>0$.
More technical details on the above decay rate estimates can be found in [11]. For the sake of clarity, we have assumed here $x\in\mathbb{R}$, however, the above result can be easily extended to the case $x\in\mathbb{R}^{n}$ (see, e.g., [9, 11]).